OpenStudy (anonymous):

The position function of a particle in rectilinear motion is given by s(t) = 2t3 – 21t2 + 60t + 3 for t ≥ 0. Find the position and acceleration of the particle at the instant the when the particle reverses direction. Include units in your answer.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8 someone plz help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it reversed direction when the derivative is zero

OpenStudy (misty1212):

find the derivative set it equal to zero and solve save that number find the second derivative plug that number in

OpenStudy (anonymous):

can u please show me the steps written out? :) to double check myself, plus it helps me learn what exact steps to follow

OpenStudy (misty1212):

did you find the derivative? it is a polynomial so that part should be easy

OpenStudy (anonymous):

yes i will do that right now :)

OpenStudy (anonymous):

6t^2-42t+60

OpenStudy (anonymous):

s(t) = 6t^2 - 42t + 60

OpenStudy (misty1212):

looks reasonable now set \[6t^2-42t+60=0\] and solve a good first step would be to divide all by 6 and solve \[t^2-7t+6=0\] which fortunately factors

OpenStudy (misty1212):

except i made a dumb mistake (blonde) it is \[t^2-7t+10=0\] which also factors

OpenStudy (anonymous):

haha its okay! thanks for the heads up :) so it would be: ( t + 5 ) ( t + 2) = 0

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

\[(t-5)(t-2)=0\] should work better

OpenStudy (anonymous):

im sorry! ahah i didn't catch that :P thank u again :)

OpenStudy (misty1212):

which is a good thing because your answers need to be positive, and now they are

OpenStudy (anonymous):

:) yes

OpenStudy (anonymous):

so i get t = 5 and t = 2? :D

OpenStudy (anonymous):

and do i plug those into s(t) =2t^3 - 21t^2 + 60t + 3

OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910 okay im all caught up to here, can u please help me with the rest of the math?!? :)

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@freckles @iambatman

OpenStudy (anonymous):

@unicwaan